OpenStudy (anonymous):

GRaph the quadratic function. Give the vertex, axis of symmetry, domain, and range. f(x)=x^2-4x+3

OpenStudy (anonymous):

@whpalmer4?

OpenStudy (ranga):

y = x^2 - 4x + 3 = x^2 - 3x - x + 3 = x(x-3) - 1(x-3) = (x-1)(x-3). The roots of this quadratic function are: 1 and 3. Also it is a parabola that opens upward and it crosses the x-axis at x = 1 and x = 3.

OpenStudy (ranga):

The x-coordinate of the vertex of the general quadratic function ax^2 + bx + c is -b/(2a). Therefore, for x^2 - 4x + 3, the x-coordinate of the vertex is: -(-4)/(2*1) = 2 y = x^2 - 4x + 3 at x = 2, y = 4 - 8 + 3 = -1 The vertex of the parabola is (2, -1)

OpenStudy (whpalmer4):

Yep, what he said! :-)

OpenStudy (whpalmer4):

Here's a graph for your viewing pleasure.

OpenStudy (ranga):